OpenStudy (anonymous):

The cost of a jacket increased from $65.00 to $78.00. What is the percentage increase of the cost of the jacket?

OpenStudy (anonymous):

83.3333333

OpenStudy (anonymous):

% change = (100%)(new-old)/old

OpenStudy (anonymous):

the answer choice is 20% 80% 13% 2%

OpenStudy (anonymous):

just make a proportion 65/78=x/100.... 65*100/78

OpenStudy (anonymous):

83.33333 is not an answer

OpenStudy (anonymous):

(100%) (78-65)/65= ?

OpenStudy (anonymous):

20%